OpenStudy (anonymous):

Mr. Arturo makes two square pans of gelatin. He cuts one pan into 39 more squares than the other pan. How many squares of gelatin does he cut in each pan?

OpenStudy (anonymous):

basically you have a problem asking you to find two numbers that are both squares that are 39 apart. To solve, you can consider the sides of each pan to be separate variables. then, you'll have the equation: a^2=b^2+39 you can then substitute different numbers in for each variable until you have two that work. you'll eventually have a = 8 b = 5 or, in the equation form: 8^2 = 5^2 +39 64= 25 + 39